OpenStudy (anonymous):

Equation containing a radical? The answer would be great, but I also want someone to show me how to solve it. Please help!!

OpenStudy (anonymous):

\[\frac{ 6 }{ \sqrt{3+2} }\]

OpenStudy (phi):

that is not an equation. but you can re-write it I would simplify 3+2 then multiply top and bottom by the sqrt(5) to "rationalize" it

OpenStudy (anonymous):

Equation could be linear or quadratic. Let's have a linear equation with a radical \[x-\sqrt 2=2\] Can you solve it?

OpenStudy (anonymous):

\[x = 2 + \sqrt{2}\] ?

OpenStudy (anonymous):

yes, you're correct.

OpenStudy (anonymous):

Okay, well then how should I simplify the one I gave?

OpenStudy (phi):

simplify 3+2 then multiply top and bottom by the sqrt(5) to "rationalize" it

OpenStudy (anonymous):

\[\frac{ 6\sqrt{5} }{ 5 }\]? But that isn't one of the multiple choice options I've been given.

OpenStudy (phi):

what are your choices ?

OpenStudy (anonymous):

A)\[\frac{ 4 }{ \sqrt{3} }\] B)\[12 - 6\sqrt{3}\] C)\[\sqrt{3-2}\] D) 3

OpenStudy (phi):

is the problem really \[ \frac{6}{\sqrt{3} +2} \] ?

OpenStudy (phi):

multiply top and bottom by the "conjugate" \[ \sqrt{3}-2 \] this is like multiplying (a+b)(a-b)= a^2 - b^2 which gets rid of the radical in the bottom (but moves it to the top)

OpenStudy (anonymous):

Oh my goodness, I didn't see that! Sorry to have wasted your time, and thank you for pointing out that mistake to me. Au revoir